OpenStudy (anonymous):

What is the slope of the line? y=2x-7

OpenStudy (anonymous):

Up 2 over 1 :D

OpenStudy (anonymous):

?

OpenStudy (anonymous):

2/1 is the slope

OpenStudy (anonymous):

there arent fractions as answers... :/

OpenStudy (anonymous):

What are the answers?

OpenStudy (anonymous):

Its a box where i type in. I cant put fractions in it.

OpenStudy (anonymous):

Oh in that case it's just 2

OpenStudy (anonymous):

oh, okay. cx Could you help me with 4 more?

OpenStudy (anonymous):

I can try!

OpenStudy (anonymous):

One sec...

OpenStudy (anonymous):

What is the y-intercept of the line? y=-7x+1

OpenStudy (anonymous):

y intercept=____________

OpenStudy (anonymous):

It's 1

OpenStudy (anonymous):

What is the slope, and what is the y-intercept of the line?

OpenStudy (anonymous):

\[y=\frac{ 2 }{ 3 } x-4\]

OpenStudy (anonymous):

Slope is -7. It's in y=mx+b form.

OpenStudy (anonymous):

Oh i thought you were talking about the top one xD one sec!

OpenStudy (anonymous):

The slope is a fraction, the y intercept isnt

OpenStudy (anonymous):

The y intercept is -4 as it is again in y=mx+b form and the slope is 2/3. Let me explain a bit to see if it's easier y=mx+b form is y=(slope)x+(Y intercept)

OpenStudy (anonymous):

For the last 2 questions, would you like me to create another question/post thing so I can give you another medal?

OpenStudy (anonymous):

Nah it's fine. Medals rant as important to me as helping :D

OpenStudy (anonymous):

*arnt

OpenStudy (anonymous):

Aweh. c: thats nice.

OpenStudy (anonymous):

What is the slope-intercept form of this equation? 6x + 4y = 8

OpenStudy (anonymous):

a. y=- 3/2 (<--- thats a fraction) x+4

OpenStudy (anonymous):

b. y= 3/2 x +4

OpenStudy (anonymous):

This is the y=mx+b form that i've been mentioning. Our slope for this is -3/2 and our y intercept is 2. Soooo y=-3/2+2

OpenStudy (anonymous):

Oh, thats what c was. cx

OpenStudy (anonymous):

Theres an X after the -3/2 xD

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

yes. cx Okay, this is the last one, but it has a graph. Im going to put the link in a reply... one sec

OpenStudy (anonymous):

Oh jesus this is a threatening graph O.O

OpenStudy (anonymous):

Which of the lines graphed in the diagram represents the equation 6x + 4y = 8? A. line c B. line b C. line a D. line d

OpenStudy (anonymous):

Isn't it?

OpenStudy (anonymous):

Okay our slope is -3/2 and our y intercept is 2. The only line fitting this requirement is line D

OpenStudy (anonymous):

True. I think I understand a little more about intercepts and slopes. Thank you

OpenStudy (anonymous):

100% c: thank you!

OpenStudy (anonymous):

np! Good luck! Let me know if I can help again!!

OpenStudy (anonymous):

:D CONGRATS!